# Trade-off of Security, Latency, and Throughput of the Nakamoto Consensus

### v3

@@ Abstract.
This paper delves into the fundamental trade-off between security, latency, and throughput in proof-of-work longest-chain-fork-choice protocols, also known as the Nakamoto consensus. New upper and lower bounds on the probability of violating transaction safety are derived as a function of honest and adversarial mining rates, an upper bound on block propagation delays, and transaction confirmation latency, both in time and in block depth. The results include a first closed-form finite-latency bound applicable to all delays and mining rates up to the ultimate fault tolerance. Notably, the gap between the upper and lower bounds is significantly narrower than the best gaps previously established for Bitcoin and its derivatives such as Litecoin and Dogecoin, as well as for Ethereum Classic. Furthermore, the paper reveals a fundamental trade-off between transaction throughput and confirmation latency, ultimately determined by the desired fault tolerance and the growth of block propagation delay as block size increases.
The authors are with the Department of Electrical and Computer Engineering, Northwestern University, Evanston, Illinois. Email: {shujie.cao, dguo}@northwestern.edu.
copyright: none
1. Introduction
Nakamoto consensus, the celebrated consensus mechanism used in Bitcoin, centers around proof-of-work (PoW) mining and the longest-chain-fork-choice rule. In particular, a transaction can be committed once it is confirmed by a sufficiently long chain or after a sufficient amount of time. For example, a widely adopted rule of thumb is to accept a Bitcoin transaction after six confirmations.
The safety of a transaction may be violated due to forks caused by block propagation delays and/or adversarial mining. Suppose the honest miners collectively mine $h$ blocks per second and the adversary mines $a$ blocks per second. Suppose also once a block is known to one honest miner, it is known to all honest miners within $\Delta$ seconds. In 2020, Dembo et al. (dembo2020everything, ) and Gaži et al. (gazi2020tight, ) used very different methods to determine the ultimate asymptotic fault tolerance for the Nakamoto consensus. Specifically, as long as [^1] The condition is expressed as $\beta<(1-\beta)/(1+(1-\beta)\lambda\Delta)$ in (dembo2020everything, ), where $\lambda=a+h$ denotes the total mining rate and $\beta=a/(a+h)$ denotes the fraction of adversarial mining.
(1) $\frac{1}{a}>\Delta+\frac{1}{h},$
every transaction is safe with arbitrarily high probabilityafter sufficient confirmation; conversely, if $1/a<\Delta+1/h$, no transaction is ever safe. How safe is a transaction after a given confirmation latency? Gaži, Ren, and Russell (gazi2022practical, ) provided a safety guarantee via numerically evaluating a Markov chain, which is only practical for small confirmation latencies and parameters that are significantly lower than the fault tolerance. By analyzing races between renewal processes, Li, Guo, and Ren (li2021close, ) presented a closed-form result focusing on confirmation by time. Using a reduction technique, Guo and Ren (guo2022bitcoin, ) gave a simple closed-form latency–security guarantee focusing on confirmation by depth, which is numerically effective across a wide range of parameters, although there remains a gap between their fault tolerance and (1). In this paper, we put forth a continuous-time model for block mining processes, blockchains, and safety violation. We develop a new closed-form latency–security trade-off for the Nakamoto consensus which is the first to guarantee safety for all parameter values of $a$, $h$, and $\Delta$ within the ultimate fault tolerance (1). The new results here also substantially reduce the previous gap between upper and lower bounds on the latency-security trade-off in many scenarios of interest.
adversary new lower new upper (guo2022bitcoin, ) upper (gazi2022practical, ) upper
10% 6 7 7 7
20% 13 14 17 15
30% 32 35 42 N/A
40% 122 161 187 N/A
49.2% 26720 52807 $\infty$ N/A
@@ 1. Introduction
Table 1. Bounds on confirmation depth for 99.9% safety. $a+h=1/600$ (blocks/second); $\Delta=10$ seconds.
In Table 1 we compare our new bounds to the best prior bounds on the minimum confirmation depth that guarantees 99.9% safety. The total mining rate is assumed to be Bitcoin’s 1 block every 600 seconds. The propagation delay bound $\Delta$ is set to 10 seconds. Our results encompass various adversarial mining fractions from $10\%$ to $49.2\%$. We include bounds for 10% and 20% adversarial mining taken directly from (gazi2022practical, ), which provides no result for higher adversarial mining rates because the numerical method therein is only feasible for parameters that are significantly distant from the ultimate fault tolerance. Evidently, the new upper bound is much closer to the lower bound than prior bounds. Moreover, the new bounds apply to all parameter values satisfying (1), which in this case is up to $a/(a+h)=49.8\%$. In contrast, the fault tolerance in (guo2022bitcoin, ) is $49.16\%$, hence for $49.16\%$, it is indicated by $\infty$ in Table 1. The inverse of the right hand side of (1) is known to be the minimum growth rate of honest blocks’ heights (pass2017analysis, ; dembo2020everything, ; gazi2020tight, ). If this growth rate exceeds the adversarial mining rate (as in (1)), the chance that more adversarial blocks are mined to overwhelm honest blocks to create a violation is very small for a confirmation latency sufficiently high. However, even if fewer adversarial blocks are mined than the height growth of honest blocks (i.e., the adversary has a deficit), a violation is still possible when honest miners with different views extend different chains to balance each other’s blocks on enough heights (i.e., the deficit is covered by those “balanced heights”). Examining the number of balanced heights offers a novel approach to the latency-security problem. Upon each opportunity to have a balanced height, there is a chance that all honest miners fail to balance, rather they synchronize to a common highest chain. Since honest blocks’ height growth rate exceeds the adversarial mining rate, there is a good chance that the adversary is left behind permanently, so no alternative highest chain is ever produced to cause a violation of the target transaction. Using this insight, we show that the number of balanced heights is stochastically dominated by essentially a geometric random variable. Interestingly, the aforementioned synchronizing block fall under the category of “Nakamoto blocks” defined in (dembo2020everything, ). Without using the complicated machinery of trees and tines as in (dembo2020everything, ; gazi2022practical, ), our analysis yields finite-latency, easily computable bounds that are tighter than the best existing results. In particular, the new bounds extend to the ultimate fault tolerance established in (dembo2020everything, ; gazi2022practical, ). The tight latency-security trade-off enables a user to choose an efficient confirmation latency that aligns with their desired security level. There has been ample evidence (e.g., (fechner2022calibrating, )) that the block propagation delay is approximately an affine function of the block size. In this paper, we explore the trade-off between the total throughput in kilobytes (KB) per second and the confirmation latency (in seconds), both of which depend on the block size and mining rate. This paper provides improved guidance for selecting parameters that optimize key performance indicators. The remainder of this paper is organized as follows: Sec. 2, introduces a mathematical model for the Nakamoto consensus. Sec. 3 explains why the model is sufficient for our purposes. Sec. 4 presents the main theorems. Sec. 5 introduces the most essential elements of the proof, followed by a sketch of the analysis. Sec. 6 presents the numerical results, showcasing the practical implications. Due to space limitations, detailed proofs are relegated to the appendices.
2. Model
@@
In this section, we put forth a mathematical model for block mining processes and blockchains, enabling a precise definition of key concepts such as confirmation, latency, and safety violations. In Sec. 3, we will delve into why the model is sufficient for the latency-security analysis.
2.1. Block and Chain
@@
We consider a continuous-time model, where transactions can be created and blocks can be mined at arbitrary times. Throughout this paper, “by time $t$” means “during $(0,t]$”.
@@
The genesis block, also referred to as block 0, is mined first. Subsequent blocks are numbered sequentially in the order they are mined, starting from block 1, block 2, and so on. From block 1 onward, each block has a unique parent block, which is mined strictly before it. The parent of block 1 is the genesis block. We let $t_{b}$ denote the time when block $b$ is mined.
Definition 2 (transaction).
Transactions are created sequentially in time as transaction 1, transaction 2, and so on. Block $b$ is associated with an array $B_{b}$, which contains the indexes of some transactions that are created strictly before time $t_{b}$. We say transaction $j$ is included in block $b$ if $j$ is in the array $B_{b}$. A transaction can be included in multiple blocks.
@@
A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$, block $b_{i-1}$ is block $b_{i}$’s parent for every $i=1,\dots,n$, and no transaction is included in two or more of those blocks. Since every chain can be uniquely identified by its last block, we also refer to the preceding chain as chain $b_{n}$. As a convention, we say the height of this chain is $n$. In general, we use $h_{b}$ to denote the height of chain $b$, which is also referred to as the height of block $b$. Evidently, the genesis block is the only block on height 00.
@@
An example of transactions, blocks, and chains.
Figure 1. An example of transactions, blocks, and chains. Nonpublic blocks are shaded. Each adversarial block is marked by a pair of horns on top. Below each block is an array of transactions included in the block.
Fig. 1 illustrates an example of chains formed by blocks 0 to 10 along with a time axis. It depicts a snapshot at $t_{10}$ when block $10$ is mined. Each block is horizontally positioned at its mining time and is connected by a line to its unique parent to its left and to its children to its right (if any). Under each block is its associated array describing the transactions included in the block. Chain $2$’s height is $h_{2}=2$; chain $5$’s height is $h_{5}=3$. We say chain $2$ is lower than chain $5$. Even though block $6$ is mined after block $5$, chain 5 is higher than chain $6$. Block $7$ or chain $7$ is the highest at its mining time. Block $1$ is associated with $B_{1}=[2,5]$, i.e., it includes transactions $2$ and $5$. Transaction $4$ is included in both blocks 2 and 3.
We next define two crucial attributes of transactions, blocks, and chains.
Definition 4 (public).
A transaction, block, or chain may become public at a specific point in timestrictly after its creation and will remain public at all times afterward. A chain being public at $t$ is equivalent to all of its blocks being public at $t$. While all transactions included in a public block are public, a transaction that is public at time $t$ may or may not be included in any block mined by $t$. As a convention, the genesis block becomes public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$.
Every chain consists of a public prefix chain followed by some nonpublic blocks (possibly none). In Fig. 1, the snapshot at $t_{10}$, we shade blocks which are nonpublic. Chain 5 is the public prefix of chain 9. If chain 9 becomes public at a later time, then block 7 must also be public by that time. Transaction 8 may be nonpublic at $t_{10}$, while transaction 11 is public since it is included in public block 6.
@@
A chain is said to be credible at time $t$, or $t$-credible, if it is no shorter than any public chain at $t$. All blocks in a credible chain are said to be credible.
Being public or not is an inherent attribute of a chain, while credibility is a different attribute determined by the set of public chains. Either public or not, a chain is $t$-credible if it is no shorter than any public chain at time $t$. In the snapshot at $t_{10}$ depicted in Fig. 1, chains 4 and 5 are both the highest public chains, making them both $t_{10}$-credible. Chain 6 is lower than chain 5, so it is not $t_{10}$-credible. Chains 7, 8, 9 and 10 are all $t_{10}$-credible. If chain 9 becomes public at a later time, then chain 8 will no longer be credible, regardless of whether it is public at that time.
2.2. Confirmation and Safety
Definition 6 (Confirmation by time $t$).
Let the confirmation time be fixed and denoted by $t>0$. Suppose a transaction becomes public at time $s>0$. We say the transaction is confirmed (by time $t$) by chain $c$ at time $s+r$ if chain $c$ is public and credible at $(s+r)$ with $r\geq t$ and contains a block which includes the transaction.
In Fig. 1, fix confirmation time $t$. Target transaction $4$ becomes public at $s$. Transaction $4$ (in block 2) is confirmed by chain 8 at $t_{10}$, which is $t_{8}$-credible, with $t_{8}>s+t$.
Definition 7 (confirmation by depth $k$).
Let the confirmation depth be a fixed natural number denoted by $k$. Suppose a transaction becomes public at $s>0$. Let block $d$ be the highest public block at time $s+\Delta$. If there exists a credible public chain $c$, which contains a block $b$ that includes the transaction, where $h_{c}\geq h_{b}+k-1$ and $h_{c}\geq h_{d}+k$, then we say chain $c$ confirms the transaction by depth $k$.
Fix $k=3$. In Fig. 1, transaction 4 is included in block 2 ($h_{2}=2$), which is contained in chain 8 ($h_{8}=4$). The highest public block by $s$ is block 1 with $h_{1}=1$ (assuming block 2 is not yet public at $s$). Since chain 8 is public and credible at $t_{10}$, $h_{8}\geq h_{2}+k-1$, and $h_{8}\geq h_{2}+k$, we say transaction 4 is confirmed (by depth $k$) by chain 8.
With confirmation by either time or depth, we can define the violation of a target transaction’s safety as follows:
Definition 8 (safety violation of a target transaction).
Suppose a target transaction becomes public at $s>0$. If there exist $r$ and $r^{\prime}$ with $r^{\prime}\geq r$ and also:
  • a chain $c$ that is public and credible at time $r$ which confirms the transaction in block $b$, and
  • an $r^{\prime}$-credible chain which does not contain block $b$,
then we say the target transaction’s safety is violated.
In Fig. 1, target transaction $4$’s safety is violated as it is confirmed by $t_{10}$-credible and public chain 8 (in block 2), while $t_{10}$-credible chain 7 does not contain block 2. Note that if two credible chains confirm the same target transaction but in different blocks, its safety is still violated.
2.3. Adversary and Propagation Delay

### v5

# Security, Latency, and Throughput of Proof-of-Work Nakamoto Consensus

@@ Abstract
This paper investigates the fundamental trade-offs between block safety, confirmation latency, and transaction throughput of proof-of-work (PoW) longest-chain fork-choice protocols, also known as PoW Nakamoto consensus. New upper and lower bounds are derived for the probability of block safety violations as a function of honest and adversarial mining rates, a block propagation delay limit, and confirmation latency measured in both time and block depth. The results include the first non-trivial closed-form finite-latency bound applicable across all delays and mining rates up to the ultimate fault tolerance. Notably, the gap between these upper and lower bounds is narrower than previously established bounds for a wide range of parameters relevant to Bitcoin and its derivatives, including Litecoin and Dogecoin, as well as Ethereum Classic. Additionally, the study uncovers a fundamental trade-off between transaction throughput and confirmation latency, ultimately determined by the desired fault tolerance and the rate at which block propagation delay increases with block size.
I Introduction
Nakamoto consensus is a foundational protocol in blockchain technology, widely adopted across cryptocurrencies to enable decentralized agreement on the blockchain’s state. In these systems, participants organize incoming transactions into blocks, with each block pointing to a unique preceding block, thereby forming a linked chain of blocks. Bitcoin and some of its derivatives use a proof-of-work (PoW) Nakamoto consensus mechanism, where miners continually compete to solve complex cryptographic puzzles; solving a puzzle grants a miner the right to add a new block to the system. The reader is referred to [1, 2] for detailed descriptions of the consensus mechanism.
A fork occurs when two or more blocks point to the same parent block. Forks can arise from adversarial mining or naturally due to block propagation delays, leading honest miners to have differing views of the blockchain state. Nakamoto consensus is built around the longest-chain fork-choice rule, which mandates that an honest miner: 1) always mines at the tip of a longest chain in its view, and 2) immediately shares any new chain extensions with other miners through a peer-to-peer network.
The most widely adopted rule is to commit a block once it is confirmed by a sufficient number of subsequent blocks. A rule of thumb for Bitcoin is six block confirmations. As an alternative, a block on a longest chain may be committed if a sufficient amount of time has lapsed since the block was seen. In either scenario, the safety of a committed block could still be violated if another chain, which excludes that block, is to emerge as a longest chain at any subsequent time. Since such a violation presents a double-spending opportunity with potentially grave consequences, it is crucial to make the probability of violation very small under the adversary’s strongest feasible attacks.
[…]
$\frac{1}{a}>\Delta+\frac{1}{h},$ (1)
every block is safe with arbitrarily high probability after sufficient block confirmations, regardless of the adversary’s attack strategy; conversely, if $1/a<\Delta+1/h$, no block is ever safe in the sense that a well-known attack can violate its safety with probability 1, regardless of the number of block confirmations.
If (1) answers the “capacity” question, then a “finite-blocklength” question ensues: Suppose the parameters satisfy (1), how safe is a committed block after $k$ block confirmations? Gaži, Ren, and Russell [5] provided a safety guarantee via numerically evaluating a Markov chain, which is only practical for small confirmation depths and parameters that are far away from the ultimate fault tolerance. By analyzing races between renewal processes, Li, Guo, and Ren [6] presented a closed-form result focusing on confirmation by time. Guo and Ren [7] introduced a rigged model, where honest blocks mined within $\Delta$ seconds of any previous blocks are flipped to become adversarial blocks. With this additional advantage, the adversary’s optimal attack reduces to the well-understood private mining attack. A simple closed-form latency-security guarantee is given in [7], although the rigging technique also weakens the fault tolerance from (1) to $\beta<1-(1/2)e^{\lambda\Delta}$, where $\lambda=a+h$ denotes the total mining rate and $\beta=a/(a+h)$ denotes the fraction of adversarial mining. Doger and Ulukus [8] refined the rigged model of [7] by not flipping honest blocks which see no other honest blocks within the $\Delta$ period before its mining, thereby improving the security-latency bounds. The same authors generalized their study based on the rigged model to the cases of different network delay assumptions in [9] and [10]. However, all of the preceding rigged models transfer some honest mining rate to the adversary, yielding strictly positive losses in the fault tolerance, i.e., (1) is not achieved.
In this paper, we carefully define a continuous-time model for block mining processes, blockchains, and safety violation for blocks. We develop a new closed-form latency-security trade-off for the Nakamoto consensus which is the first to achieve the ultimate fault tolerance (1). The new results here also substantially reduce the previous gap between upper and lower bounds on the latency-security trade-off in many scenarios of interest.
$\beta$ new lower new upper [7] [8] [5]
10% 6 8 7 7 7
20% 13 16 17 15 15
30% 32 39 42 38 N/A
40% 122 156 187 165 N/A
49.2% 26720 44024 $\infty$ $\infty$ N/A
@@ I Introduction
TABLE I: Bounds on confirmation depth for 99.9% safety. $\lambda=1/600$ (blocks/second); $\Delta=10$ seconds.
In Table I we compare our new bounds to the best prior bounds on the minimum confirmation depth that guarantees 99.9% safety. The total mining rate is assumed to be Bitcoin’s 1 block every 600 seconds. The propagation delay bound $\Delta$ is set to 10 seconds. Our results encompass various adversarial mining fractions from $10\%$ to $49.2\%$. We include bounds for 10% and 20% adversarial mining taken directly from [5], which provides no result for higher adversarial mining rates because the numerical method therein is infeasible for parameters that are close to the fault tolerance. Evidently, the new upper bound is much closer to the lower bound than prior bounds in challenging cases. Moreover, the new bounds apply to all parameter values satisfying (1), which in this case is up to $\beta=49.8\%$. In contrast, the fault tolerance in [7] is $49.16\%$, hence for $49.2\%$, it is indicated by $\infty$ in Table I.
The inverse of the right-hand side of (1) is known to be the minimum growth rate of honest blocks’ heights [11, 3, 4]. If this growth rate exceeds the adversarial mining rate (as in (1)), the chance that more adversarial blocks are mined to overwhelm honest blocks to create a violation is very small for a confirmation latency sufficiently high. However, even if fewer adversarial blocks are mined than the height growth of honest blocks (i.e., the adversary has a deficit), a violation is still possible when honest miners with different views extend different chains to balance each other’s blocks on enough heights (i.e., the deficit is covered by those “balanced heights”). This paper develops a novel approach to the latency-security problem by examining the number of balanced heights.
Upon each opportunity to have a balanced height, there is a chance that all honest miners fail to balance, rather they synchronize to a common highest chain. Since honest blocks’ height growth rate exceeds the adversarial mining rate, there is a good chance that the adversary is left behind permanently, so no alternative highest chain is ever produced to cause a violation of the target block. Using this insight, we show that the number of balanced heights is stochastically dominated by essentially a geometric random variable. Interestingly, the aforementioned synchronizing block falls under the category of “Nakamoto blocks” defined in [3]. Without using the complicated machinery of trees and tines as in [3, 5], our analysis yields tight, easily computable finite-latency security bounds which achieve and imply the ultimate fault tolerance (1).
The tight latency-security trade-off enables a user to choose an efficient confirmation latency that aligns with their desired security level. There has been ample evidence (e.g., [12]) that the block propagation delay is approximately an affine function of the block size. In this paper, we explore the trade-off between the total throughput in kilobytes (KB) per second and the confirmation latency (in minutes), both of which depend on the block size and mining rate. This paper provides improved guidance for selecting parameters that optimize key performance indicators.
The remainder of this paper is organized as follows: Sec. II, introduces a mathematical model for the Nakamoto consensus. Sec. III explains why the model is sufficient for our purposes. Sec. IV presents the main theorems. Sec. V introduces the most essential elements of the proof, followed by a sketch of the analysis. Sec. VI presents the numerical results, showcasing the practical implications. Detailed proofs are relegated to the appendices.
II Model
@@
In this section, we put forth a mathematical model for block mining processes and blockchains, enabling a precise definition of key concepts such as confirmation, latency, and safety violations. In Sec. III, we will delve into why the model is sufficient for the latency-security analysis.
II-A Block and Chain
@@
We consider a continuous-time model, where blocks can be mined at arbitrary times. Throughout this paper, “by time $t$” means “during $(0,t]$”.
@@
The genesis block, also referred to as block 0, is mined first. Subsequent blocks are numbered sequentially in the order they are mined, starting from block 1, block 2, and so on. From block 1 onward, each block has a unique parent block, which is mined strictly before it. The parent of block 1 is the genesis block. We use $t_{b}$ to denote the time when block $b$ is mined.
@@
A sequence of $n+1$ block numbers, $(b_{0},b_{1},\dots,b_{n})$, defines a blockchain (or simply a chain) if $b_{0}=0$, and for every $i=1,\dots,n$, block $b_{i-1}$ is block $b_{i}$’s parent. As a convention, we say the height of this chain is $n$. Since every chain can be uniquely identified by its last block, we often refer to the preceding chain as chain $b_{n}$. Also, we say block $b_{i+1}$ extends chain $b_{i}$ and that chain $b_{i}$ is a prefix of chain $b_{n}$ for every $i=0,\dots,n-1$. In general, we use $h_{b}$ to denote the height of chain $b$, which is also referred to as the height of block $b$. Evidently, the genesis block is the only block on height 00.
@@
Figure 1: An example of blocks and chains. Nonpublic blocks are shaded. Each adversarial block is marked by a pair of horns on top.
Fig. 1 illustrates an example of chains formed by blocks 0 to 10 along with a time axis. It depicts a snapshot right after $t_{10}$ when block $10$ is mined. Each block is horizontally positioned at its mining time and is connected by a line to its unique parent to its left and to its children to its right (if any). Chain $2$’s height is $h_{2}=2$; chain $5$’s height is $h_{5}=3$. We say chain $2$ is lower than chain $5$. Even though block $6$ is mined after block $5$, chain 5 is higher than chain $6$. Blocks $9$ and $10$ or chains $9$ and $10$ are both the highest at $t_{10}$.
Definition 3 (public and public height).
A chain may become public at a specific time strictly after its creation and will remain public at all times afterward. If a chain is public, all of its prefix chains must also be public. As a convention, the genesis block is public at time 0 and block 1 is mined afterward, i.e., $t_{1}>0$. At time $t$, the height of a highest public chain is referred to as the public height.
We use the notion of being public to capture the situation that a chain is known to all honest nodes. Every chain consists of a public prefix chain followed by some nonpublic blocks (possibly none). In Fig. 1, the snapshot at $t_{10}$, we shade blocks which are nonpublic. Chain 5 is the public prefix of chain 9. If chain 9 becomes public at a later time, then chain 7 must also be public by that time.
@@
A chain mined by time $t$ is said to be credible at $t$, or $t$-credible, if it is no lower than the public height at $t$.
Being public or not is an inherent attribute of a chain, while credibility is a different attribute determined by the set of public chains. Either public or not, a chain is $t$-credible if it is no shorter than any public chain at time $t$. In the snapshot at $t_{10}$ depicted in Fig. 1, chain 8 is the highest public chain. Chains 4, 5, 6 are lower than chain 8, so they are not $t_{10}$-credible. Chains 7, 8, 9 and 10 are all $t_{10}$-credible. If chain 10 becomes public at a later time, then chain 7 will no longer be credible, regardless of whether it is public at that time.
II-B Confirmation and Safety
Definition 5 (confirmation by time).
Let the confirmation time be fixed and denoted by $t>0$. Given a time $s$, if a chain $g$ is mined by $s$ and is a prefix of a $(s+r)$-credible chain $c$ with $r\geq t$, then we say chain $c$ confirms chain $g$ by time of $t$.
An example is depicted in Fig. 1 with $s$ and $s+t$ marked on the time axis. Suppose chain 1 is mined by time $s$. Since chain 7 is credible at $t_{7}>s+t$, it confirms chain 1 by time of $t$. Likewise, chain 8 confirms chain 2 by time of $t$.
Definition 6 (confirmation by depth).
Let the confirmation depth be a fixed natural number denoted by $k$. If a block $g$ is included in a credible chain $b$ with $h_{b}\geq h_{g}+k-1$, then we say chain $b$ confirms block $g$ by depth of $k$ or block $g$ is $k$-confirmed by chain $b$.
In Fig. 1, for example, block 2 is 3-confirmed by chain 8, and block 3 is 3-confirmed by chain 7 and 4-confirmed by chain 9. With confirmation by either time or depth, we can define the violation of a block’s safety as follows:
Definition 7 (safety violation of a block).
If there exist $r$ and $r^{\prime}$ with $r^{\prime}\geq r$ as well as
  • an r-credible chain $c$ which confirms block $g$, and
  • an $r^{\prime}$-credible chain which does not include block $g$,
then we say block $g$’s safety is violated.
In Fig. 1, block 3 is 3-confirmed by $t_{7}$-credible chain 7, but is not included in chain 8, which is $t_{8}$-credible. Thus block 3’s safety is violated.
II-C Adversary and Propagation Delay